OpenStudy (anonymous):

x^-3 = 1/27 How do you get x?

OpenStudy (johnweldon1993):

how else can you write x^-3? 1/x^3 right? looking any better?

OpenStudy (johnweldon1993):

so now just ask yourself....what ^3 is 27?

OpenStudy (anonymous):

cross multiply

OpenStudy (mertsj):

\[\frac{1}{x^3}=\frac{1}{27}\]

OpenStudy (mertsj):

Cross multiply

OpenStudy (anonymous):

1/3^3 = 1/27 is that right?

OpenStudy (johnweldon1993):

correct....x = 3

OpenStudy (anonymous):

ok thanks